OpenStudy (anonymous):

can someone explain to me what angular frequency is? and how it relates to period, frequency and amplitude?

OpenStudy (unklerhaukus):

Consider a particle moving in uniform circular motion on the trajectory of a unit circle. The time it takes to all the way around, is the period, say for example, the period T = 1 [s] The frequency is the inverse of the period, i.e. f = 1 / T , and in our case = 1 [1/s] = 1 [Hz] (one time per second == one hertz) The angular frequency is like the frequency, but is in terms of radians per second angular frequency: ω = 2πf, which makes = 2π [rad/second] Which just mean the particle moves through 2π radians (a complete revolution) every second.

OpenStudy (unklerhaukus):

i suppose the amplitude of either x, or y coordinate of the position of the particle, would be 1, because we chose a unit circle as the trajectory, ..

OpenStudy (irishboy123):

\.

OpenStudy (anonymous):

I finally get It!! thank you!!

OpenStudy (unklerhaukus):

\[\ddot\smile\]